Jasmine has $4.90 in quarters, dimes, nickels, and pennies. She has 2 more dimes than quarters. She has 3 more nickels than quarters. She has 4 more pennies than quarters. How many quarters does Jasmine have? Note: A quarter is worth $0.25, a dime is worth $0.10, a nickel is worth $0.05, and a penny is worth $0.01.
Answer by mananth(16949)


Jasmine has $4.90 in quarters, dimes, nickels, and pennies.
let Jasmine have x quarters
She has 2 more dimes than quarters.
d= x+2
She has 3 more nickels than quarters.
n=x+3
She has 4 more pennies than quarters.
p=x+4
Jasmine has $4.90 = 490 cents
25x+10(x+2)+5(x+3)+x+4 = 490
25x +10x+20+5x+15 +x+4=490
41x = 451
x= 11
11 quarters